The x-coordinate of the incentre of the triangle that the coordinates of mid points of its sides as (0,1) (1,1)and (1,0) is

Grade:11

1 Answers

SHAIK AASIF AHAMED
9 years ago
let A(x1,y1) B(x2,y2) C(x3,y3) are the vertices of a triangle then
x1+x2=0; x2+x3=2; x3+x1=2
y1+y2=2; y2+y3=2; y3+y1=0
solving above equations we get A(0,0) B(0,2) C(2,0)
now a=BC=2.828 ,b=CA=2 ,c=AB=2
incentre x coordinate is given by\frac{ax1+bx2+cx3}{a+b+c}
which is 2-sqrt(2)=0.585
Thanks and Regards
Shaik Aasif
askIITians Faculty